# On the Real Zeroes of Half-integral Weight Hecke Cusp Forms, II

### v1

@@ 10. A random computation
Our next goal is to estimate the contribution of primes of different size.
@@ 10.1. Contribution of the primes with $K^{\theta_{0}}<p\leq K^{\theta_{J}}$
$=\sum_{p|m_{1}m_{2}\Rightarrow p\in I_{j}}\frac{\chi_{d}(m_{1}m_{2})}{m_{1}^{s_{1}+1/2}m_{2}^{s_{2}+1/2}}X(m_{1})X(m_{2})\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{j}\\ \Omega(n)\leq 4\ell_{j}\end{subarray}}\frac{\lambda(n)\chi_{d}(n)\nu_{2}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)X(u).$
@@
$\sum_{p|m_{1}m_{2}\Rightarrow p\in I_{j}}\frac{\chi_{d}(m_{1}m_{2})}{m_{1}^{s_{1}+1/2}m_{2}^{s_{2}+1/2}}\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{j}\\ \Omega(n)\leq 4\ell_{j}\end{subarray}}\frac{\lambda(n)\chi_{d}(n)\nu_{2}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)\sum_{h|(m_{1},m_{2})}1_{m_{1}m_{2}/h^{2}=u}$
(10.2) $=\sum_{(h,d)=1}\frac{1}{h^{1+s_{1}+s_{2}}}\sum_{p|m_{1}m_{2}\Rightarrow p\in I_{j}}\frac{\chi_{d}(m_{1}m_{2})}{m_{1}^{s_{1}+1/2}m_{2}^{s_{2}+1/2}}\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{j}\\ \Omega(n)\leq 4\ell_{j}\\ m_{1}m_{2}|n\end{subarray}}\frac{\lambda(n)\chi_{d}(n)\nu_{2}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}c_{n}(m_{1}m_{2}).$
@@
$\Sigma(z):=\sum_{(h,d)=1}\frac{1}{h^{1+s_{1}+s_{2}}}\sum_{p|m_{1}m_{2}\Rightarrow p\in I_{j}}\frac{\chi_{d}(m_{1}m_{2})}{m_{1}^{s_{1}+1/2}m_{2}^{s_{2}+1/2}}\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{j}\\ m_{1}m_{2}|n\end{subarray}}\frac{z^{\Omega(n)}\lambda(n)\chi_{d}(n)\nu_{2}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}c_{n}(m_{1}m_{2}).$
Now this expression can be written as an Euler product
$\Sigma(z)=\prod_{\begin{subarray}{c}p\in I_{j}\\ p\nmid d\end{subarray}}\left(1-\frac{z}{\sqrt{p}}\left(\frac{1}{p^{s_{1}+1/2}}+\frac{1}{p^{s_{2}+1/2}}\right)+\frac{z^{2}}{2p}\left(\frac{1}{p^{1+2s_{1}}}+\frac{1}{p^{1+2s_{2}}}+\frac{1}{p^{1+s_{1}+s_{2}}}\right)+O\left(\frac{1}{p^{3/2+2\min\{\textnormal{Re}(s_{1}),\textnormal{Re}(s_{2})\}}}\right)\right)$
From this it is easily seen that $|\Sigma(z)|\ll 1$ uniformly for $|z|\leq 2$ when $\min\{\textnormal{Re}(s_{1}),\textnormal{Re}(s_{2})\}>1/2$. Applying this bound in (10.3) we see that
@@
10.2. The contribution of the small primes
Next we need to understand the contribution of the primes with $c\leq p\leq K^{\theta_{0}}$. This involves understanding the interaction between $L(s_{1}+1/2,s_{2}+1/2;X,d)$ and $M_{2,0}(X;d)$. A key point is that since $M_{2,0}(X;d)$ consists of relatively small primes we can express it in terms of an Euler product with negligible loss since $\ell_{0}$ is large. This allows us to simplify our later analysis by reducing the problem to understanding the contribution from each prime $p\in I_{0}$ individually. Let
@@ 10.2. The contribution of the small primes
$R(X;d):=\widetilde{M}_{2,0}(X;d)-M_{2,0}(X;d)=\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{0}\\ \Omega(n)>4\ell_{0}\end{subarray}}\frac{\lambda(n)\chi_{d}(n)\nu_{2}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)X(u).$
Let us denote $L_{0}(s_{1},s_{2};X,d):=\prod_{p\in I_{0}}L_{p}(s_{1},s_{2};X,d)$. Then by the Cauchy–Schwarz inequality we have
@@
$\sum_{\begin{subarray}{c}p|n_{1}\Rightarrow p\in I_{0}\\ \Omega(n_{1})>4\ell_{0}\end{subarray}}\sum_{\begin{subarray}{c}p|n_{2}\Rightarrow p\in I_{0}\\ \Omega(n_{2})>4\ell_{0}\end{subarray}}\frac{\lambda(n_{1})\lambda(n_{2})\chi_{d}(n_{1}n_{2})\nu_{2}(n_{1};\ell_{0})\nu_{2}(n_{2};\ell_{0})}{2^{\Omega(n_{1})+\Omega(n_{2})}\sqrt{n_{1}n_{2}}}\sum_{u_{1}|n_{1}}\sum_{u_{2}|n_{2}}c_{n_{1}}(u_{1})c_{n_{2}}(u_{2})\mathbb{E}(X(u_{1})X(u_{2})).$
@@
$\mathbb{E}(|R(X;d)|^{2})\leq\frac{1}{2^{8\ell_{0}}}\sum_{p|r\Rightarrow p\in I_{0}}\frac{C^{\Omega(r)}}{r}\leq\frac{(\log K)^{O(1)}}{2^{4\ell_{0}}},$
where we have used the fact that $c$ is sufficiently large so that the sum in the middle converges. Combining the two estimates above completes the proof. ∎
10.3. Estimates for primes $p>K^{\theta_{J}}$
@@
For $\textnormal{Re}(s_{1}),\,\textnormal{Re}(s_{2})>1/2$ we have that
@@
$=1_{p\nmid d}\cdot\zeta_{p}(1+s_{1}+s_{2})$
for $\textnormal{Re}(s_{1}),\,\textnormal{Re}(s_{2})>1/2$. ∎
Thus using the dominated convergence theorem we have for $\textnormal{Re}(s_{1}),\,\textnormal{Re}(s_{2})>1/2$ and any $z\geq 2$ that
@@
We further compute
@@
$=2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\theta_{J}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)d(u)}{\sqrt{u}}\chi_{d}(nu)\left(\log\left(\frac{Kd}{2\pi\sqrt{u}}+O(1)\right)\right)$
$-2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\theta_{J}}}\frac{h_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\varepsilon)}e^{2s_{1}^{2}}u^{-s_{1}}v^{2s_{1}}\,\frac{\mathrm{d}s_{1}}{s_{1}^{2}}+O\left(KX(\log K)^{-7}\right).$
@@
$=4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\mathbb{E}\left(\frac{1}{(2\pi i)^{2}}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}\left(\frac{2\pi}{|d|}\right)^{-s_{1}-s_{2}}\frac{\Gamma(s_{1}+k)}{\Gamma(k)}\cdot\frac{\Gamma(s_{2}+k)}{\Gamma(k)}e^{s_{1}^{2}+s_{2}^{2}}F(s_{1},s_{2};X,d)\frac{\mathrm{d}s_{1}\,\mathrm{d}s_{2}}{s_{1}s_{2}}\right),$
@@
Recall that $\mathbb{E}(F(s_{1},s_{2};X,d))$ has a pole at $s_{1}+s_{2}=0$.
We move the line of integration from $\textnormal{Re}(s_{2})=\varepsilon$ to $\textnormal{Re}(s_{2})=-1/2$, crossing simple poles at $s_{2}=0$ and $s_{2}=-s_{1}$. The $s_{2}$-integral is $O(K^{-1/2+\varepsilon}X^{1/2})$ on the new line. Next we shall study the contribution of the residues at these poles to (10.8) separately.
@@
(10.9) $4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\frac{1}{2\pi i}\int\limits_{(\varepsilon)}\left(\frac{2\pi}{|d|}\right)^{-s_{1}}\frac{\Gamma(s_{1}+k)}{\Gamma(k)}e^{s_{1}^{2}}\zeta(1+s_{1})\prod_{p\leq x}\zeta_{p}(1+s_{1})^{-1}\prod_{\begin{subarray}{c}p>x\\ p|d\end{subarray}}\zeta_{p}(1+s_{1})^{-1}$
@@
in the region $\textnormal{Re}(s_{1})\geq\frac{1}{2}+\varepsilon$. Since for fixed $\sigma>0$ Stirling’s formula gives $|\Gamma(\sigma+it)|\ll(|t|+1)^{\sigma-\frac{1}{2}}e^{-\pi|t|}$, by (10.10) we may truncate the $s_{1}$-integral (10.9) to $|\textnormal{Im}(s_{1})|\leq B\log K$ at the cost of an error term of size $O(K^{-1})$, where $B$ is a sufficiently large absolute constant.
@@
By a straightforward computation using the residue theorem and (10.3) shows that the double pole contributes the amount
$4(\log K)\sum_{n\leq K^{4\theta_{J}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)\chi_{d}(nu)}{\sqrt{u}}$
@@
to the (10.9). Here $\gamma$ is the Euler–Mascheroni constant and
@@
$4(\log K)\sum_{n\leq K^{4\theta_{J}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)d(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\left(\log\left(\frac{kd}{2\pi\sqrt{u}}\right)+\sum_{p|d}\frac{\log p}{p-1}+\gamma\right).$
@@
$2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\theta_{J}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)d(u)}{\sqrt{u}}\chi_{d}(nu)\left(\log\left(\frac{Kd}{2\pi\sqrt{u}}+O(1)\right)\right).$
@@
$-4(\log K)\sum_{n\leq K^{4\theta_{J}}}\frac{h_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{d}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)$
$\qquad\qquad\times\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\varepsilon)}\frac{\Gamma(s_{1}+k)}{\Gamma(k)}\cdot\frac{\Gamma(-s_{1}+k)}{\Gamma(k)}e^{2s_{1}^{2}}u^{-s_{1}}v^{2s_{1}}\,\frac{\mathrm{d}s_{1}}{s_{1}^{2}}.$
@@
(10.11) $-2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\theta_{J}}}\frac{h_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\varepsilon)}e^{2s_{1}^{2}}u^{-s_{1}}v^{2s_{1}}\,\frac{\mathrm{d}s_{1}}{s_{1}^{2}}+O\left(K^{3/4+\varepsilon}X^{3/4}\right).$
@@
$=2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\theta_{J}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)d(u)}{\sqrt{u}}\chi_{d}(nu)\left(\log\left(\frac{Kd}{2\pi\sqrt{u}}+O(1)\right)\right)+O\left(X^{3/4}K^{1+\varepsilon}\right).$

### v2

@@ 10. A random computation
Our next goal is to estimate the contribution of primes of different sizes.
@@ 10.1. Contribution of the primes with $K^{\theta_{0}}<p\leq K^{\theta_{J}}$
$=\sum_{p|m_{1}m_{2}\Rightarrow p\in I_{j}}\frac{\chi_{d}(m_{1}m_{2})}{m_{1}^{s_{1}+1/2}m_{2}^{s_{2}+1/2}}X(m_{1})X(m_{2})\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{j}\\ \Omega(n)\leq 4\ell_{j}\end{subarray}}\frac{\lambda(n)\chi_{d}(n)\nu_{4}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)X(u).$
@@
$\sum_{p|m_{1}m_{2}\Rightarrow p\in I_{j}}\frac{\chi_{d}(m_{1}m_{2})}{m_{1}^{s_{1}+1/2}m_{2}^{s_{2}+1/2}}\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{j}\\ \Omega(n)\leq 4\ell_{j}\end{subarray}}\frac{\lambda(n)\chi_{d}(n)\nu_{4}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)\sum_{h|(m_{1},m_{2})}1_{m_{1}m_{2}/h^{2}=u}$
(10.2) $=\sum_{(h,d)=1}\frac{1}{h^{1+s_{1}+s_{2}}}\sum_{p|m_{1}m_{2}\Rightarrow p\in I_{j}}\frac{\chi_{d}(m_{1}m_{2})}{m_{1}^{s_{1}+1/2}m_{2}^{s_{2}+1/2}}\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{j}\\ \Omega(n)\leq 4\ell_{j}\\ m_{1}m_{2}|n\end{subarray}}\frac{\lambda(n)\chi_{d}(n)\nu_{4}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}c_{n}(m_{1}m_{2}).$
@@
$\Sigma(z):=\sum_{(h,d)=1}\frac{1}{h^{1+s_{1}+s_{2}}}\sum_{p|m_{1}m_{2}\Rightarrow p\in I_{j}}\frac{\chi_{d}(m_{1}m_{2})}{m_{1}^{s_{1}+1/2}m_{2}^{s_{2}+1/2}}\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{j}\\ m_{1}m_{2}|n\end{subarray}}\frac{z^{\Omega(n)}\lambda(n)\chi_{d}(n)\nu_{4}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}c_{n}(m_{1}m_{2}).$
Changing the order of summation, the inner double sum can be written as an Euler product
$\sum_{p|n\Rightarrow p\in I_{j}}\frac{z^{\Omega(n)}\lambda(n)\chi_{d}(n)\nu_{4}(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{\begin{subarray}{c}m_{1},m_{2}\\ m_{1}m_{2}|n\end{subarray}}\frac{\chi_{d}(m_{1}m_{2})}{m_{1}^{s_{1}+1/2}m_{2}^{s_{2}+1/2}}c_{n}(m_{1}m_{2})$
$=\prod_{\begin{subarray}{c}p\in I_{j}\\ p\nmid d\end{subarray}}\left(1-\frac{z}{\sqrt{p}}\left(\frac{1}{p^{s_{1}+1/2}}+\frac{1}{p^{s_{2}+1/2}}\right)+\frac{3z^{2}}{2p}\left(1+\frac{1}{p^{1+2s_{1}}}+\frac{1}{p^{1+2s_{2}}}+\frac{1}{p^{1+s_{1}+s_{2}}}\right)+O\left(\frac{1}{p^{3/2+\min\{\textnormal{Re}(s_{1}),\textnormal{Re}(s_{2})\}}}\right)\right)$
From this it is easily seen that $|\Sigma(z)|\ll 1$ uniformly for $|z|\leq 2$ when $\min\{\textnormal{Re}(s_{1}),\textnormal{Re}(s_{2})\}>0$ (recall that $j\neq 0$). Applying this bound in (10.3) we see that
@@
10.2. Contribution of the small primes
Next we need to understand contribution of the primes with $c\leq p\leq K^{\theta_{0}}$. This involves understanding the interaction between $L(s_{1}+1/2,s_{2}+1/2;X,d)$ and $M_{2,0}(X;d)$. A key point is that since $M_{2,0}(X;d)$ consists of relatively small primes we can express it in terms of an Euler product with negligible loss since $\ell_{0}$ is large. This allows us to simplify our later analysis by reducing the problem to understanding the contribution from each prime $p\in I_{0}$ individually. Let
@@ 10.2. Contribution of the small primes
$R(X;d):=\widetilde{M}_{2,0}(X;d)-M_{2,0}(X;d)=\sum_{\begin{subarray}{c}p|n\Rightarrow p\in I_{0}\\ \Omega(n)>4\ell_{0}\end{subarray}}\frac{\lambda(n)\chi_{d}(n)\nu_{2}(n;\ell_{j})}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)X(u)$
and let us denote $L_{0}(s_{1},s_{2};X,d):=\prod_{p\in I_{0}}L_{p}(s_{1},s_{2};X,d)$. Then by the Cauchy–Schwarz inequality we have
@@
$\sum_{\begin{subarray}{c}p|n_{1}\Rightarrow p\in I_{0}\\ \Omega(n_{1})>4\ell_{0}\end{subarray}}\sum_{\begin{subarray}{c}p|n_{2}\Rightarrow p\in I_{0}\\ \Omega(n_{2})>4\ell_{0}\end{subarray}}\frac{\lambda(n_{1})\lambda(n_{2})\chi_{d}(n_{1}n_{2})\nu_{4}(n_{1};\ell_{0})\nu_{4}(n_{2};\ell_{0})}{2^{\Omega(n_{1})+\Omega(n_{2})}\sqrt{n_{1}n_{2}}}\sum_{u_{1}|n_{1}}\sum_{u_{2}|n_{2}}c_{n_{1}}(u_{1})c_{n_{2}}(u_{2})\mathbb{E}(X(u_{1})X(u_{2})).$
@@
$\mathbb{E}(|R(X;d)|^{2})$ $\leq\frac{1}{2^{8\ell_{0}}}\sum_{p|r\Rightarrow p\in I_{0}}\sum_{p|n_{1}\Rightarrow p\in I_{0}}\sum_{p|n_{2}\Rightarrow p\in I_{0}}\frac{\lambda(rn_{1})\lambda(rn_{2})\chi_{d}(r^{2}n_{1}n_{2})\nu(rn_{1})\nu(rn_{2})}{r\sqrt{n_{1}n_{2}}}c_{rn_{1}}(r)c_{rn_{2}}(r)$
$\leq\frac{(\log K)^{O(1)}}{2^{8\ell_{0}}}\sum_{p|r\Rightarrow p\in I_{0}}\frac{C^{\Omega(r)}}{r}$
(10.6) $\leq\frac{(\log K)^{O(1)}}{2^{8\ell_{0}}},$
where we have used the fact that $c$ is sufficiently large so that the sum in the middle converges. Combining the two estimates above and recalling that $\eta_{1}>0$ in the definition of $\ell_{0}$ is taken to be sufficiently large completes the proof. ∎
10.3. Estimates for the large primes
@@
For $\min\{\textnormal{Re}(s_{1}),\textnormal{Re}(s_{2})\}>1/2$ we have that
@@
$=1_{p\nmid d}\cdot\zeta_{p}(1+s_{1}+s_{2}).$
Thus using the dominated convergence theorem we have for $\min\{\textnormal{Re}(s_{1}),\textnormal{Re}(s_{2})\}>1/2$ and any $z\geq 2$ that
@@
We further compute using the Hecke relations for $X(n)$ and (6.2) that
@@
$=2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\delta_{0}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)d(u)}{\sqrt{u}}\chi_{d}(nu)\left(\log\left(\frac{Kd}{2\pi\sqrt{u}}\right)+O(1)\right)$
$-2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\delta_{0}}}\frac{h_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\frac{1}{2}+\varepsilon)}e^{2s_{1}^{2}}u^{-s_{1}}v^{2s_{1}}\,\frac{\mathrm{d}s_{1}}{s_{1}^{2}}$
$\qquad\qquad\qquad\qquad\qquad\qquad+O\left(KX(\log K)^{-7}\right).$
@@
(10.10) $=4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)$
$\qquad\qquad\times\mathbb{E}\left(\frac{1}{(2\pi i)^{2}}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(\frac{1}{2}+\varepsilon)}\left(\frac{2\pi}{|d|}\right)^{-s_{1}-s_{2}}\frac{\Gamma(s_{1}+k)}{\Gamma(k)}\cdot\frac{\Gamma(s_{2}+k)}{\Gamma(k)}e^{s_{1}^{2}+s_{2}^{2}}F(s_{1},s_{2};X,d)\frac{\mathrm{d}s_{1}\,\mathrm{d}s_{2}}{s_{1}s_{2}}\right),$
@@
Recall that $\mathbb{E}(F(s_{1},s_{2};X,d))$ has a simple pole at $s_{1}+s_{2}=0$.
We move the line of integration from $\textnormal{Re}(s_{2})=1/2+\varepsilon$ to $\textnormal{Re}(s_{2})=-1$, crossing simple poles at $s_{2}=0$ and $s_{2}=-s_{1}$. The $s_{2}$-integral is $O(K^{-1/2+\varepsilon}X^{1/2})$ on the new line. Next we shall study the contribution of the residues at these poles to (10.9) separately.
@@
(10.11) $4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\frac{1}{2\pi i}\int\limits_{(\frac{1}{2}+\varepsilon)}\left(\frac{2\pi}{|d|}\right)^{-s_{1}}\frac{\Gamma(s_{1}+k)}{\Gamma(k)}e^{s_{1}^{2}}\zeta(1+s_{1})\prod_{p\leq K^{\theta_{J}}}\zeta_{p}(1+s_{1})^{-1}\prod_{\begin{subarray}{c}p>K^{\theta_{J}}\\ p|d\end{subarray}}\zeta_{p}(1+s_{1})^{-1}$
@@
in the region $\textnormal{Re}(s_{1})\geq\frac{1}{2}+\varepsilon$. Since for fixed $\sigma>0$ Stirling’s formula gives $|\Gamma(\sigma+it)|\ll(|t|+1)^{\sigma-\frac{1}{2}}e^{-\pi|t|}$, by (10.12) we may truncate the $s_{1}$-integral (10.11) to $|\textnormal{Im}(s_{1})|\leq B\log K$ at the cost of an error term of size $O(X)$, where $B$ is a sufficiently large absolute constant.
@@
A straightforward computation, using the residue theorem and (10.3), shows that the double pole contributes the amount
$4(\log K)\sum_{n\leq K^{4\delta_{0}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)\chi_{d}(nu)}{\sqrt{u}}$
@@
to (10.11). Here $\gamma$ is the Euler–Mascheroni constant
@@
$4(\log K)\sum_{n\leq K^{4\delta_{0}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)d(u)}{\sqrt{u}}\chi_{d}(nu)$
$\qquad\qquad\qquad\times\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\left(\log\left(\frac{k|d|}{2\pi\sqrt{u}}\right)+\sum_{p|d}\frac{\log p}{p-1}+\gamma\right).$
@@
$2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\theta_{J}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)d(u)}{\sqrt{u}}\chi_{d}(nu)\left(\log\left(\frac{K|d|}{2\pi\sqrt{u}}\right)+O(1)\right).$
@@
$-4(\log K)\sum_{n\leq K^{4\delta_{0}}}\frac{h_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)$
$\qquad\qquad\times\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\frac{1}{2}+\varepsilon)}\frac{\Gamma(s_{1}+k)}{\Gamma(k)}\cdot\frac{\Gamma(-s_{1}+k)}{\Gamma(k)}e^{2s_{1}^{2}}u^{-s_{1}}v^{2s_{1}}\,\frac{\mathrm{d}s_{1}}{s_{1}^{2}}.$
@@
(10.13) $-2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\delta_{0}}}\frac{h_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\frac{1}{2}+\varepsilon)}e^{2s_{1}^{2}}u^{-s_{1}}v^{2s_{1}}\,\frac{\mathrm{d}s_{1}}{s_{1}^{2}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+O\left(K^{3/4+\varepsilon}X^{3/4}\right).$
@@
$=2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\delta_{0}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)d(u)}{\sqrt{u}}\chi_{d}(nu)\left(\log\left(\frac{K|d|}{2\pi\sqrt{u}}\right)+O(1)\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad+O\left(X^{3/4}K^{1+\varepsilon}\right).$